I need help to solve the equation 2x - 3y = -7 using slope and y-intercept

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

I need help to solve the equation 2x - 3y = -7 using slope and y-intercept
2x-3y= -7
-3y= -7-2x
Divide the equation by -1
3y= 7+2x
or
3y = 2x+7
y= 2/3 * x +7/2
this is the form of equation y= mx+c m the slope and c the intercept

Answer by ikleyn(53996)

.

How the problem is worded in the post, this formulation is mathematically incorrect.

A correct formulation is THIS

    Transform this given equation to the slope-intercept form.